Of first-time college students who matriculate to a certain university, the odds in favor of having graduated in the top 25 percent of their high school class are 2.65 to 1. Of transfer students who matriculate to the same university, 0.5 proportion graduated in the top 25 percent of their high school class.

(a) For first-time college students, the proportion who graduated in the top 25 percent of their high school class is _____ (rounded to three decimal places)

(b) For transfer students, the odds in favor of having graduated in the top 25 percent of their high school class are _____ to 1.
Answer #1

solution:

a)proportion=odds/(1+odds)

=2.65/1+2.65

=2.65/3.65

=0.72602

=0.726(round to three decimals)

b)odds=proporiton/(1-proportion)

=0.5/1-0.5

=0.5/0.5

  =1

classes are 1 to 1